A particle is fired vertically upward with a speed of 15km/s. With what speed will it in interstellar space. Assume only earth's gravitational field.

Grade:Upto college level

1 Answers

Pratham Ashish
14 years ago

hi,

 to be free from earth,s field & enter in interstellar space one must have a minimum energy corresponding to the escape velocity.

escape velocity of earth = 11.2 km/s  = 11.2  X 103 m / s

      corresponding  K.E. = 1/2 m (11.2  X 103)^2

let the remaining speed in  interstellar space is v km/sec

  remaining K.E.

                     1/2 m (vx 103)^2 =  1/2 m ( 15 X 103 ) ^2 - 1/2 m ( 11.2 x 103 ) ^2

                              v^2 =  15 ^2 - 11.2 ^2

                                     = 99.56

                               v = √99.56 km/ sec